OpenStudy (anonymous):

Calculate the dissociation energy of NaBr(g). Express your answer in kJ/mol. Data NaBr Madalung constant 1.75 NaBr Born exponent 8 Na+ Radius 116e-12 m Br+ Radius 182e-12 m K+ Radius 152e-12 m

OpenStudy (anonymous):

408 KJ/mole

OpenStudy (anonymous):

how did you get that???

OpenStudy (anonymous):

E = e^2x(4xπxε0xr0)^-1x(1-n^-1); e- = 1.602*10^-19 eV; ε0 = 8.854*10^-12; r0 = (116+182) = 298*10^-12 m; E = 6.77*10-19 J or 408 KJ/mol.

OpenStudy (anonymous):

cool guys

OpenStudy (anonymous):

where did u guys keep madalungs constant

OpenStudy (anonymous):

madalungs constant for crystals or solid this is a gas

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what if you are calculating the cohesive energy

OpenStudy (anonymous):

E= -M*NA*e2*1/(4*π*ε0*r0)*(1-1/n); e-= 1.602*10^-19; ε0= 8.854*10^-12; r0= (116+182): 298*10^-12; M= 1.75; n= 8. E = - M*e^2x(1/4xπxε0xr0)x(1-1/n).

OpenStudy (anonymous):

i did all that i seem not to pay attention to my calculations very much i had the wrong answer but i am fine with it Thanks anyway